OpenStudy (wasiqss):

Integrate {(1+x)/(1-x)}^1/2

OpenStudy (wasiqss):

wth is that

OpenStudy (wasiqss):

k.

OpenStudy (anonymous):

Hm you guys need me ?

OpenStudy (wasiqss):

yehh!!

OpenStudy (anonymous):

I have solved it.

OpenStudy (wasiqss):

wow ffm u r genious :)

OpenStudy (anonymous):

put x = cos(t)

OpenStudy (katrinakaif):

Wait ^ Solving

OpenStudy (wasiqss):

im waiting :P

OpenStudy (anonymous):

Put \( x = \cos 2z\), if I am not making any mistake your equation would get \(-4 \int \cos^2 z dz \)

OpenStudy (anonymous):

and then it's easy.

OpenStudy (mimi_x3):

thats all? wtf

OpenStudy (anonymous):

@sam:I find sub-multiples difficult to handle while integrating.

OpenStudy (wasiqss):

ffm is it cox^2z or cos2z

OpenStudy (anonymous):

Alright guys time to give medals :P

OpenStudy (anonymous):

I have used latex so I believe it's quite easy to comprehend no ?

OpenStudy (anonymous):

@formath \[-2\sqrt{2}\int\limits_{}\cos ^{3}{} z dz\]

OpenStudy (anonymous):

and Mimi, your answer is not a valid step.

OpenStudy (mimi_x3):

why isn't it valid?

OpenStudy (wasiqss):